Download the annual report of any commercial bank for year 2018 and 2019 and

perform liquidity and profitability analysis and interpret the result


1
Expert's answer
2021-06-09T12:26:18-0400

A case of Bahrain

(a) liquidity analysis

Current ratio=current assets÷current liabilities

2018

="\frac{12.1}{57.9}=0.209"

2019

"\frac{15.4}{62.3}=0.247"

The current ratio increased from 2018 to 2019 which means that the firm is able to meets it's short term obligations.


(b) profitability analysis

=sum of customer paid-(amortized fixed cost+variable cost)

2018

"70.2-(17.7+12.3)=40.2"

2019

"90.7-(22.1+10)=58.6"

The firm is thriving well as it's making profits.